OpenStudy (anonymous):

A funnel with a height of 3.67 inches is shown below. The funnel has a diameter of 1.87 inches that includes a wall that is 0.125 inches thick. What is the maximum volume of the funnel rounded to the nearest hundredth?

OpenStudy (campbell_st):

find the volume of the inner funnel radius = (0.935 -0.125) I don't have a picture so I'll assume its con shaped \[V = 1/3 \pi r^2 h\] \[V = 1/3 \times \pi \times 0.815^2 \times3.67\]